# Understanding the Challenges and Promises of Developing Generative AI Apps: An Empirical Study

### v3

@@ 5.2. RQ2 Findings
7) Comparison to Other Apps. Comparison to Other Apps refers to comparing the AI app to other tools users have previously tried.
Figure 14. Examples of reviews discussing Comparisons to Other Apps
Insights. We find that 66% of the reviews draw comparisons to other AI-powered apps (e.g., comparing Microsoft Bing Search with Google). We also observe that 9% of the reviews compare the app with non-Gen-AI alternatives. For instance, review 1 in Figure 14 criticizes the forced integration of the AI chat feature, expressing a strong preference for traditional search functionality. Finally, 18% of reviews draw comparisons between the apps they review and general-purpose AI tools (e.g., ChatGPT), highlighting expectations shaped by user experiences with AI tools, as shown in review 2 in Figure 14. 77% of the reviews that compare the app with ChatGPT believe that ChatGPT outperforms the app in question, while 15% prefer the performance of the domain-specific app. These comparisons highlight a key challenge for domain-specific Gen-AI app developers. Their apps must offer distinct value that goes beyond the general-purpose functionality of ChatGPT.
8) Customer Support & Community. Customer Support & Community refers to reviews that address developers directly to express their experience, such as reporting user enjoyment with specific features or requesting to fix app shortcomings.
Figure 15. Examples of reviews discussing Customer Support & Community
Insights. In our sample, we find that 51% of users express enjoyment with specific features (e.g., review 1 in Figure 15), 19% suggest new features to enhance their experience (e.g., review 2 in Figure 15), 9% request to fix app shortcomings (e.g., review 3 in Figure 15), and 21% discuss generic enjoyment of the app without relating it to a specific feature.
$\Rightarrow$ Implication 4.
ChatGPT is a baseline comparator to users, in order to succeed developers must offer value beyond what ChatGPT offers. Developers should also consider creating a toggle button to allow users to opt in or out of AI-enabled features, such as AI search. The regular features may be sufficient for users (e.g., regular search can satisfy users’ searching needs, and AI search is not always necessary). User reviews highlight key features that developers should preserve when updating their apps. Prior work provided approaches that facilitate competitor analysis (Maram_LLM_Cure; PeerApps_SANER; Assi_Hassan_Tian_Zou_2021). For competing developers, analyzing user feedback from similar apps can also uncover desirable features worth adopting to remain competitive in the market.
9) Human AI Interaction & Emotional Connection. Human AI Interaction & Emotional Connection reviews highlight the emotional and relational dimensions of human–AI interaction, particularly with conversational Gen-AI apps.
@@
Figure 16. Examples of reviews discussing Human AI Interaction & Emotional Connection
Insights. We observe that users enjoy their interactions with conversational Gen-AI apps for multiple reasons. First, users enjoy a realistic interaction (i.e., closely similar to human dialogue). This perceived realism fosters a sense of companionship and emotional connection (e.g., review 1 in Figure 16). Second, reviews highlight the comfort in the emotional support provided by Gen-AI apps, particularly because the interaction lacks real human judgment, as shown in review 2 in Figure 16. This illustrates the importance of AI’s perceived neutrality and non-judgmental nature, especially when discussing sensitive or emotionally charged topics. Finally, users point to the therapeutic potential of conversational AI in providing mental and emotional support, such as managing mental health challenges (e.g., review 3 in Figure 16) and developing strategies to achieve personal goals.
10) Accessibility & Inclusivity. Accessibility and inclusivity refer to the degree to which the Gen-AI apps accommodate users for affordability, multilingual support, age appropriateness, and usability for individuals with cognitive or physical impairments.
Figure 17. Examples of reviews discussing Accessibility & Inclusivity
Insights. Users discuss different accessibility and inclusivity topics including: 1) financial barriers (raised in 31% of the reviews, such as review 1 in Figure 17), 2) greater language support (raised in 22% of the reviews, such as review 2 in Figure 17), 3) disability-related needs (raised in 8% of the reviews, such as reviews 3 and 4 in Figure 17), 4) age-related concerns (raised in 10% of the reviews, such as Reviews 5 and 6 in Figure 17), and 5) general accessibility concerns (raised in 31% of the reviews).
$\Rightarrow$ Implication 5.
The perception of conversational Gen-AI apps as emotionally supportive and non-judgmental suggests their growing role as trusted companions, particularly in contexts involving emotional vulnerability. Developers should account for these evolving user expectations by ensuring that Gen-AI interfaces maintain a tone of empathy and neutrality, while also incorporating safeguards to avoid over-reliance or the illusion of professional psychological care. Furthermore, this trend highlights the need for interdisciplinary collaboration with mental health professionals to responsibly shape the affective design of Gen-AI systems used for emotionally sensitive interactions.
The findings further highlight the importance of designing Gen-AI apps with diverse accessibility needs in mind. Developers should prioritize inclusive design by integrating features such as read-aloud support, customizable text presentation, and adaptive interfaces tailored to neurodiverse users (e.g., Attention Deficit Hyperactivity Disorder, Dyslexia and Dyspraxia).
Observation 2.4: Developer Responses to Gen-AI Reviews Are Primarily Courteous or Diagnostic, With Few Immediate Resolutions. Table 8 shows that for Gen-AI apps, the most common developer response type is Acknowledgment and Appreciation (31%), which consists largely of brief thank-you messages. These replies most often accompany positive reviews and serve as courteous engagement rather than substantive technical follow-up.
The second most frequent category, Request for Additional Information (30%), reflects the practical limits of troubleshooting Gen-AI issues through review threads. Even when space is not a constraint, users rarely provide the detailed context needed to diagnose problems that relate to model behavior, device configuration, or account-specific issues. Developers therefore redirect users to email or in-app support channels where logs, device information, or credentials can be shared securely.
By contrast, Issue Resolved and Update Advised appears rarely (2%), whereas Acknowledged and Under Investigation is more common (10%). This pattern aligns with the nature of Gen-AI systems: many issues require model adjustments, infrastructure changes, or updates to safety and moderation mechanisms, none of which can be addressed immediately or within the constraints of a review thread. As a result, developers are more likely to acknowledge problems and communicate that work is underway than to confirm the issue has been resolved.
@@
Hassan et al. (developer_response_safwat) analyzed developer replies across a broad set of mobile apps and identified recurring response drivers such as “thank the user”, “ask for more details”, “provide guidance”, “ask for endorsement”, “advertise other products”, wait for updates, and “justify advertisements”. Many of these practices align closely with our taxonomy. For example, their “thank the user” category maps to our “Acknowledgment and Appreciation,” their “ask for more details” category maps to our “Request for Additional Information,” and their “justify the advertisements” category maps onto our broader Clarification or Policy Rebuttal, which includes explanations about subscription models, premium features, and content restrictions. Notably, their work does not treat “issue resolved” as a dominant response driver, and we similarly observe that confirmed resolutions are relatively rare in our Gen-AI dataset (2%). We report the category explicitly for completeness of Gen-AI developer behaviors, but its low prevalence mirrors prior findings rather than contrasting with them. Overall, the strong correspondence between the two taxonomies suggests that general response practices extend to Gen-AI apps, while our domain-specific categories capture the additional emphasis developers place on clarifying AI-related behavior and managing expectations around the evolving nature of Gen-AI features.
Observation 2.5: App categories show highly distinctive developer response patterns. Table 9 reveals variation in how developers across different app categories respond to user feedback. Some categories exhibit exceptionally high engagement, while others show very limited interaction. For example, Games stands out with both the highest reply ratio (56%) and the longest average response delay (44.7 days). In contrast, Education demonstrates the shortest response times, with developers replying in less than a day on average (0.7 days). Meanwhile, Entertainment apps show the lowest reply ratio among categories with meaningful activity (2%). These contrasting patterns suggest that developer responsiveness is shaped not only by category size but also by the operational demands, user expectations, and support norms unique to each domain.
@@
Our manual analysis of developer responses in the Games category helps explain why it exhibits both the highest reply ratio and the longest response delay. Responses in which developers report that an issue has been fixed account for 30% of all replies in this category, far exceeding the 2% observed in the overall dataset. This emphasis on confirmed fixes aligns with the long response delay shown in Table 9. This likely reflects that developers often wait until a patch is deployed before responding, resulting in slower but more substantive communication. Responses indicating that an issue is acknowledged and under investigation represent 13% of replies, slightly higher than the overall 10%. Together, these categories show that Games developers place strong emphasis on technical issue resolution and ongoing investigation.
Generic thank-you messages appear in only 12% of replies, considerably lower than the overall 31%. The remaining responses highlight active engagement with user issues: 28% request additional information, 10% provide troubleshooting guidance, and 8% offer clarifications or policy rebuttals. Overall, developer communication in Games is markedly problem-oriented and resolution-driven, which leads to both high engagement and extended response timelines.
A closer look at Education apps helps explain why this category exhibits the fastest average developer response time. In our manual sample, the majority of Education responses fall under Acknowledgment and Appreciation (73%), which is markedly higher than both the overall rate of 31% and the 12% observed in Games. These replies are typically short, template-based thank-you messages that require minimal effort to generate, allowing developers to respond quickly and consistently.
Education apps also show a comparatively high rate of Request for Additional Information (16%). This proportion is notably larger than the overall sample and reflects a common pattern in Gen-AI Education reviews. Users frequently report that the app produced incorrect answers, yet they often omit the original prompt, the model’s response, or screenshots of the interaction. Without these details, developers cannot diagnose or reproduce the issue within the review thread. As a result, they quickly reply by directing users to contact support channels where complete context can be shared securely. This response type is therefore both common and fast, since it does not require problem investigation within the review platform itself.
By contrast, response types requiring deeper engagement occur far less frequently in Education, including Guidance and Troubleshooting Tips (2%), Acknowledged and Under Investigation (6%), and Issue Resolved and Update Advised (1%). Together, these patterns show that Education apps tend to rely primarily on quick acknowledgments or requests for additional information, which contributes to the category’s exceptionally short response delay.
@@
Entertainment apps show the lowest reply ratio in our dataset, and the distribution of response types helps explain this pattern. Acknowledgment and Appreciation responses account for only 14%, far below the overall rate of 31%. This indicates that developers in this category do not typically reply to routine or positive reviews and instead reserve responses for situations where user action or developer follow-up is required. Correspondingly, a large share of replies fall under Request for Additional Information (38%) and Guidance and Troubleshooting Tips (22%), suggesting that developers primarily intervene when users report crashes, playback issues, or account problems that require diagnostic detail. The relatively high proportion of Acknowledged and Under Investigation (19%) further shows that many of these reports involve technical issues that cannot be resolved immediately, which may discourage broader engagement with lower-priority feedback. Together, these patterns indicate that Entertainment developers focus their limited response efforts on actionable problem reports rather than general user engagement, resulting in a lower overall reply ratio.
Observation 2.6: Developer responsiveness varies substantially across Gen-AI topic categories. Table 10 summarizes developer reply behavior for each Gen-AI topic category. The topic with the highest reply ratio is Accessibility & Inclusivity (47%), despite having the smallest number of reviews. This indicates that developers treat accessibility-related feedback as a high-impact area that warrants immediate and consistent attention. Such reviews often report barriers to use for specific user populations, making them critical from both ethical and product-quality perspectives. The short average delay (0.2 days) further reflects the urgency developers assign to this category.
Beyond accessibility, the next highest reply ratios are found in Utility & Use Cases (18%) and Content Quality (14%), suggesting that developers are highly responsive to feedback about how users apply AI features and whether the generated outputs meet expectations. These topics directly affect user satisfaction and daily usability, which likely motivates more frequent developer engagement.
In contrast, the categories with the lowest reply ratios are Customer Support & Community (1%), Human–AI Interaction & Emotional Connection (2%), and Content Policy & Censorship (5%). These topics often involve complaints about communication quality, emotional tone, safety constraints, or moderation decisions. Such issues are typically systematic, tied to model behavior or policy rules that cannot be quickly adjusted, which may explain the lower response frequency and the longer average delays seen in these categories.
This contrast between highly actionable categories and structurally constrained categories reinforces the idea that responsiveness is shaped by whether an issue is immediately fixable or requires deeper system changes. Overall, these patterns show that developers prioritize issues that are actionable in the short term (e.g., accessibility barriers, inaccurate outputs, task-level failures) while responding less frequently to concerns that require deeper model changes or policy revisions. This reflects a practical triaging strategy driven by the nature of Gen-AI systems and the feasibility of immediate fixes.
@@
Human–AI Interaction & Emotional Connection 25,142 542 2% 7.5

### v4

@@ 5.2. RQ2 Findings
Comparison to Other Apps D: Comparison to Other Apps captures instances where users evaluate the Gen-AI app by explicitly comparing it with alternative tools they have previously used.
E: “This is by far the best app out there it’s better than Socratic, Chegg or any other app claiming to be able to help you with your homework. It gives you a detailed answer and it gives you why that is the answer. I love this app, I cannot believe i had just found it”
Table 6. Gen-AI topics with descriptions and example reviews.
Opportunity 3: Discussions of Gen-AI versatility highlight opportunities to design flexible systems that support evolving and user-driven use cases. User discussions demonstrate that Gen-AI systems are applied across a wide variety of domains and tasks, highlighting the versatility of generative capabilities. Evidence for this opportunity emerges across several topics, particularly AI Performance and Utility & Use Cases.
Reviews discussing AI Performance indicate that users evaluate the AI based on its ability to handle diverse tasks and domains. In this topic, users comment on aspects such as accuracy (19%), response speed (18%), conversational quality (15%), and versatility across tasks (10%). Positive sentiment toward this topic is high, with 80% of reviews expressing satisfaction. Similarly, discussions in Utility & Use Cases demonstrate that users continuously discover new ways to apply Gen-AI tools in everyday contexts, including cooking advice, fitness guidance, studying religion, learning musical instruments, programming, general problem solving, and for life advice.
These findings suggest that many Gen-AI use cases emerge organically from user experimentation rather than from predefined developer intentions. For developers, this highlights the importance of monitoring how users adapt AI systems to new tasks and designing flexible systems that can accommodate evolving user needs.
Challenge 1: Discussions of user expectations and AI limitations highlight challenges in managing expectation gaps and communicating system capabilities transparently. Despite generally positive perceptions of Gen-AI capabilities, reviews highlight frustration when AI systems fail to meet user expectations. These concerns are primarily reflected in the AI Performance and Content Quality topics.
While 80% of reviews discussing AI Performance are positive, approximately 18% of reviews report performance limitations, particularly related to poor AI memory, inaccurate responses, inconsistent conversational behavior, and slow response times. Similarly, 21% of reviews discussing Content Quality report dissatisfaction with generated outputs, often describing cases where the AI fails to follow prompts accurately or produces unexpected results. These issues become particularly pronounced when users submit complex prompts requiring multiple elements or relationships, where generated outputs may become partially incorrect or inconsistent.
These findings highlight a key challenge for developers: managing the gap between user expectations and the actual capabilities of Gen-AI systems. Developers may benefit from designing mechanisms that communicate system limitations more transparently, such as providing feedback when prompts exceed model capabilities or offering suggestions for refining prompts.
Challenge 2: Discussions of content moderation and creative freedom highlight challenges in balancing safety requirements with flexible and transparent moderation systems. Content moderation in regards to censorship and restriction of content emerges as one of the most controversial aspects of Gen-AI apps. This challenge is captured in the Content Policy & Censorship topic.
User sentiment toward moderation mechanisms is mixed. While 28% of reviews express positive sentiment toward content policies, 55% express dissatisfaction. Among reviews mentioning filtering mechanisms, 77% report frustration with the system’s behavior. The most common complaint involves overly restrictive filtering, where harmless prompts are blocked unexpectedly (29%). Additionally, 23% of users express frustration that censorship limits their ability to generate mature or violent narratives, suggesting that current moderation policies restrict creative freedom. Conversely, a smaller portion of users (6%) argue that moderation systems fail to block inappropriate content effectively.
These findings highlight the difficulty of balancing safety requirements with users’ desire for creative freedom. Users often hold conflicting expectations regarding acceptable content boundaries, particularly when engaging in creative storytelling or roleplaying scenarios. For developers and platform providers, these results highlight the importance of designing transparent and adaptable moderation systems that maintain safety standards while accommodating diverse creative use cases.
Challenge 3: Discussions of Gen-AI feature integration highlight challenges in determining when AI capabilities add meaningful value while maintaining reliability. User feedback also reveals challenges related to how Gen-AI capabilities are integrated into mobile apps. This challenge is reflected in the Features & Functionality and Comparison to Other Apps topics.
Reviews discussing Features & Functionality show that users appreciate conversational AI capabilities such as voice interaction, with 68% of reviews expressing positive sentiment. However, some users report that integrating AI features into existing systems can introduce reliability issues, particularly when AI-powered assistants fail to perform simple tasks that traditional systems previously handled effectively. Additionally, reviews in the Comparison to Other Apps topic demonstrate that users frequently evaluate Gen-AI apps relative to well-known AI tools such as ChatGPT and Photomath. These comparisons suggest that users expect new Gen-AI apps to offer capabilities that match or exceed existing systems.
Together, these findings highlight an important design challenge: not every feature benefits from the integration of Gen-AI. Developers must carefully determine when AI capabilities meaningfully enhance the user experience and when simpler solutions may be more appropriate. Successfully integrating Gen-AI therefore requires thoughtful design decisions that balance innovation, reliability, and clear value for users.
Summary of RQ 2 Our analysis shows that user discussions of Gen-AI apps center on generative capabilities and practical use cases, while negative feedback is primarily driven by broader app design issues. Reviews of Gen-AI features receive higher ratings, indicating generally positive user perceptions. Qualitative findings highlight opportunities in accessibility, creative collaboration, and diverse use cases, alongside challenges in managing user expectations, balancing content moderation with creative freedom, and determining effective integration of Gen-AI features.
6. RQ3: What temporal trends emerge in Google Play Store user feedback on Gen-AI topics?
6.1. Experiment Setup
We track the evolution of user feedback over time by conducting a two-level trend analysis. First, we conduct a high-level comparison of overall trends between Gen-AI and non-Gen-AI topics by examining how 1) average ratings and 2) the percentage of reviews evolve over time. This perspective helps highlight the increasing user attention to Gen-AI features. Second, we perform a more fine-grained, category-level analysis of Gen-AI topics to uncover detailed temporal patterns. We track the evolution of individual Gen-AI topic categories using clustering techniques and qualitative review analysis. The following steps outline how we conduct the analysis:
Step 1: data preparation. We group reviews according to their topic category $(T)$ and time period $(P)$. Specifically, for every topic category $T$, we organize reviews into yearly intervals based on their submission dates. Then, for each topic category $T$, we construct two time series across all time periods $P$: 1) $Avg_{(T,P)}$ as the average rating of reviews in category $T$ posted during period $P$, and 2) $Pr_{(T,P)}$ as the percentage of Gen-AI reviews assigned to category $T$ within each app during period $P$, computed relative to the total number of Gen-AI reviews for that app in the same period. To avoid large apps dominating the results, we compute both measures at the app level and then average them across all apps that contain reviews in that period.
Step 2: clustering methodology. To identify patterns in how Gen-AI topics evolve over time, we cluster the topics based on their trends. For each time series $Avg_{(T,P)}$ and $Pr_{(T,P)}$, we follow the three-step process listed below:
  1. 1.
    Clustering algorithm. K-Means clustering is a partitioning method that groups unlabeled data into $k$ clusters by minimizing within-cluster variance (Warren Liao, 2005). In our case, each topic category $T$ is represented across all time periods $P$, using either $Avg_{(T,P)}$ or $Pr_{(T,P)}$. These values form a time series vector for each topic, capturing how the topic evolves over time. We apply K-Means clustering to group topics with similar temporal trajectories. We select K-Means because it is efficient, flexible, and widely used for clustering in data mining tasks (Ikotun et al., 2023; Assi et al., 2025b).
  2. 2.
    Optimal number of clusters. To determine the optimal number of clusters ($k$), we use the elbow method, a widely adopted technique for evaluating clustering structure by examining the reduction in within-cluster variance as $k$ increases (Ashari et al., 2023). Through elbow plot evaluation and manual inspection of clusters, we identify $k=3$ as the point where additional clusters provide only marginal improvement for both Rating and Percentage. For $Pr_{(T,P)}$, the clusters reflect increasing, decreasing, and fluctuating patterns (Figure 6). For $Avg_{(T,P)}$, the clusters correspond to increasing, decreasing, and stable trends (Figure 7).
To statistically validate the observed differences between Gen-AI and non-Gen-AI ratings over time, we perform a paired Wilcoxon signed-rank test on the yearly average ratings. Specifically, for each time period $P$, we compute the average rating of Gen-AI reviews and the average rating of non-Gen-AI reviews and compare the paired values across all periods.
Step 3: manual trend investigation. To further examine trends in average ratings, we conduct an in-depth qualitative analysis of two representative topics. First, we analyze the opportunity AI for Accessibility and Wellbeing through the Emotional Connection topic to understand how users engage with Gen-AI emotionally over time. Second, we examine the challenge Balancing Content Moderation and Creative Freedom, through the topic Content Policy & Censorship which exhibits highly variable and often conflicting user perspectives, making it particularly informative for understanding evolving user concerns.
To interpret the identified trends, we conduct a qualitative analysis of randomly sampled reviews from three key time periods. For each topic category and time period, we draw samples at a 95% confidence level with a 10% margin of error. We manually code a total of 407 reviews to identify recurring concerns, behavioral patterns, and shifts in user expectations, providing qualitative context for the observed quantitative trends.
6.2. Findings
Finding 3.1: The proportion of reviews discussing Gen-AI topics decreases over time but remains substantial. Figure 5 illustrates the proportion of reviews discussing Gen-AI topics over time. Gen-AI related reviews represent a large share of user feedback throughout the dataset, accounting for approximately 77% of reviews in 2019 and gradually declining to around 41% by 2026. Despite this decline, Gen-AI related discussions continue to represent a substantial portion of the overall review landscape. This pattern suggests that while the relative prominence of Gen-AI topics has decreased over time, Gen-AI capabilities remain an important focus of user feedback within mobile apps.
Figure 5. Evolution of the percentage of Gen-AI reviews (left), and of the evolution of the average ratings of Gen-AI vs Non-Gen-AI topics (right) (Full dataset considered).
Finding 3.2: Gen-AI reviews receive consistently higher ratings than non-Gen-AI reviews. As shown in Figure 5, reviews associated with Gen-AI topics consistently receive higher ratings than those discussing non-Gen-AI aspects across all time periods. While non-Gen-AI ratings decrease from 3.9 to 2.3, Gen-AI ratings remain higher, generally between 4.5 and 3.9. A paired Wilcoxon signed-rank test confirms this difference ($p<0.05$), indicating that Gen-AI-related feedback is rated significantly more positively than non-Gen-AI feedback.
Finding 3.3: Topic categories cluster into fluctuating, decreasing, and increasing trends based on their percentage of reviews over time. We cluster topic categories based on the temporal evolution of their review proportions into three trends as shown in Figure 6. The fluctuating cluster includes Comparison to Other Apps, Content Policy & Censorship, and Features & Functionality. These topics vary over time in response to external developments and emerging issues. Comparison to Other Apps rises sharply following the introduction of prominent Gen-AI systems (e.g., ChatGPT in 2022) (Akhtar, 2024), as users begin benchmarking apps against well-known alternatives such as GPT-based systems or domain-specific tools like Photomath. This topic continues to fluctuate as new technologies and competitors emerge. Similarly, Content Policy & Censorship fluctuates because it is driven by user dissatisfaction; users tend to discuss it primarily when encountering moderation issues, leading to intermittent spikes (notably around 2023 and 2026). Features & Functionality shows moderate variation, reflecting ongoing but less event-driven user feedback on app capabilities.
The increasing cluster includes AI Performance, Content Quality, Emotional Connection, and Utility & Use Cases. These topics show a steady rise over time, indicating that as Gen-AI systems improve, users increasingly engage with and evaluate core system capabilities. In particular, the growth of AI Performance and Content Quality reflects improving model capabilities and rising user expectations. At the same time, the increase in Emotional Connection and Utility & Use Cases suggests that users are moving beyond novelty, increasingly discussing how they interact with AI and the diverse contexts in which they apply it. Overall, these trends indicate a shift from novelty-driven engagement to performance- and use-oriented evaluation, highlighting the maturation of user expectations as Gen-AI technologies evolve.
Finding 3.4: Emotional Connection with AI evolves from novelty to dependency and addiction concerns, while declining developer engagement with these emerging risks highlights the need for responsible design and safeguards. The trend is shown in Figure 7. We summarize the sub-topics discussed in each year in Figure 8. Emotional interaction with AI shifts over time from novelty-driven companionship to more sustained and structured forms of engagement. Early use centers on conversational companionship, but with time, users increasingly rely on AI for emotional support, social confidence, and expression. As adoption matures, interactions become more habitual, with users describing consistent reliance on AI for communication and coping. This progression introduces new risks, including over-reliance, addiction, and disengagement from real-world interactions, particularly among emotionally vulnerable users. Overall, the trend reflects a transition from exploratory use to dependency-related challenges, highlighting the need for responsible design and mechanisms that promote healthy usage boundaries. Developer responsiveness in Emotional Connection declines sharply from 55% in 2022 to below 2% after 2024, despite the emergence of more complex concerns such as addiction and emotional dependency, indicating reduced attention to longer-term user risks.
@@ 6.2. Findings
Figure 8. Subtopics of reviews discussing Emotional Connection (Representative subset considered).
Figure 9. Subtopics of reviews discussing Content Policy & Censorship (Representative subset considered).
Finding 3.5: Content moderation concerns shift from trust and privacy to censorship, accompanied by increased developer responsiveness. The trend is shown in Figure 7 and the subtopics are summarized in Figure 9. Content moderation concerns evolve from early distrust around data use and privacy toward debates about censorship, bias, and platform control. As adoption matures, users focus less on whether their data is safe and more on how content is filtered and governed. A key tension emerges between users reporting excessive moderation (restrictions on regular use) and those requesting reduced moderation for greater freedom in generating mature content. While mentions of both increase over time, requests for reduced moderation grow more strongly than calls for stricter moderation in response to unintended explicit outputs. New concerns around age-related restrictions and verification further complicate user experience and trust.
Overall, the findings point to a growing mismatch between platform governance mechanisms and evolving user expectations, highlighting the need for more transparent, flexible, and user-aware moderation strategies. Developer responses increase for content-related concerns, rising from 3% in 2019 to 20% in 2026, reflecting a focus on actionable issues like censorship. In contrast, responses to emotional connection concerns decline, leaving human-centered risks comparatively under-addressed.
Summary of RQ 3 To Our temporal analysis shows that (1) the proportion of Gen-AI discussions decreases over time but remains substantial, (2) Gen-AI reviews consistently receive higher ratings than non-Gen-AI reviews, (3) emotional interaction with AI evolves from novelty to dependency concerns, and (4) content moderation challenges shift from early trust and privacy issues to tensions around censorship, bias, and platform control.
7. Discussion
In this section, we first present a platform-comparison analysis to assess the generalizability to the App Store. Second, we present an analysis of developers’ responses to explore how they respond to Gen-AI reviews.
7.1. Platform Comparative Analysis
To assess whether insights derived from Google Play Store reviews generalize to other mobile app stores, we apply SARA to the App Store versions of apps in our dataset and compare the resulting topic distributions across platforms. The analysis focuses on the 59 apps that are available on both the Google Play Store and the App Store, enabling a controlled comparison of user feedback for the same apps across different review environments.
Table 7. Distribution of Gen-AI topic categories in App Store (Astore) and Google Play Store (GPS) reviews (full dataset considered).
Topic Category Astore Reviews GPS Reviews Astore Rating GPS Rating
Utility & Use Cases 74,089 (32.9%) 111,736 (14.3%) 4.9 4.8
AI Performance 23,846 (10.6%) 130,492 (16.7%) 3.3 4.2
Content Quality 16,477 (7.3%) 86,888 (11.1%) 3.7 3.6
Content Policy & Censorship 12,299 (5.5%) 16,934 (2.2%) 2.4 2.5
Creative Potential 11,607 (5.2%) 57,551 (7.4%) 4.9 3.6
Emotional Connection 1,071 (0.5%) 6,446 (0.8%) 3.8 4.7
Comparison to Other Apps 983 (0.4%) 1,154 (0.1%) 4.4 4.4
Features & Functionality 740 (0.3%) 18,283 (2.3%) 4.0 2.3
Inclusivity 743 (0.3%) None 2.3 None
Topic categories remain consistent across platforms. Across app categories, the overlap of topic categories between Google Play Store and App Store reviews ranges from 64% to 100% (average 81%), indicating that users discuss largely similar aspects of Gen-AI apps regardless of platform. However, the emphasis placed on these topics differs across platforms.
App Store users more frequently describe how they apply Gen-AI tools in practice, whereas Google Play Store users more often evaluate the technical capabilities of the underlying Gen-AI models. As shown in Table 7, App Store reviews are dominated by discussions of Utility & Use Cases, which account for 32.9% of Gen-AI related reviews. In contrast, this topic accounts for only 14.3% of Gen-AI discussions in Google Play Store reviews, whereas AI Performance emerges as the most frequently discussed topic (16.7%).
App Store users tend to be more critical when evaluating the reliability and quality of AI-generated outputs. Discussions of AI Performance receive a lower average rating on the App Store (AGR = 3.3) than on Google Play Store (AGR = 4.2). Differences also emerge in how users discuss the Emotional Connection of Gen-AI systems. In Google Play Store reviews, this topic often reflects positive experiences such as companionship or emotional support (AGR = 4.7), while App Store reviews more frequently frame similar experiences in terms of potential risks, including addiction or negative effects on mental health (AGR = 3.8).
App Store users may be more attentive to issues related to privacy, bias, and responsible AI behavior. Ethical concerns appear more prominently in App Store reviews. Inclusivity-related issues, such as gender and ethnicity bias in generated content, form a distinct topic category in App Store reviews but do not appear frequently enough in Google Play Store reviews to constitute a separate category. Similarly, privacy concerns captured under Content Policy & Censorship appear as a topic across two App Store app categories, while they are only discussed in one Google Play Store app category.
Summary of Platform Analysis Overall, these results suggest that while the primary dimensions of user feedback are consistent across platforms, the emphasis placed on them differs. App Store users tend to highlight practical usage scenarios and ethical considerations, whereas Google Play Store users more frequently evaluate the technical performance of Gen-AI systems. Monitoring feedback across multiple platforms, therefore, provides complementary insights into user expectations and emerging concerns surrounding Gen-AI apps.
7.2. Developer Reply Analysis
This section examines how developers engage with user feedback in Gen-AI app reviews. While previous sections focused on user perspectives, developer replies reveal how teams assess reported issues, manage user expectations, and respond to the practical constraints of diagnosing AI-related problems through public review channels. For developer responses, we conduct two qualitative analyses: first, coding a representative sample of 96 replies across all app categories to characterize general response patterns, and second, analyzing 281 replies from three focal categories (Games, Education, and Entertainment) to capture category-specific developer practices.
Developer responses are predominantly generic and template-based, with limited occurrence of substantive issue resolution. The most common developer response type, as shown in Table 8, is Acknowledgment and Appreciation (31%), representing typically short thank-you messages posted in response to positive reviews.
@@ 7.2. Developer Reply Analysis
Developers redirect users to external support channels where complete context and diagnostic information can be shared securely. The second most common category, Request for Additional Information (30%), reflects the practical limitations of diagnosing Gen-AI issues through review threads. Users report incorrect or unexpected AI outputs without providing the original prompt, generated response, or screenshots. By contrast, confirmed resolutions are rare: only 2% of replies indicate that an issue has been resolved, while 10% acknowledge that developers are investigating the problem. This pattern reflects the nature of Gen-AI systems, where many issues require model adjustments, infrastructure changes, or updates to safety mechanisms that cannot be deployed immediately. As a result, review replies often serve as an initial point of contact rather than a channel for resolving AI-related failures directly.
General developer reply practices reported in prior work extend to Gen-AI apps. Hassan et al. (Hassan et al., 2018) identified several common developer response drivers in mobile app ecosystems, including thanking users, requesting additional details, and providing guidance. Many of these patterns align closely with our taxonomy. For example, their “thank the user” category corresponds to our Acknowledgment and Appreciation. Overall, this correspondence suggests that general developer reply practices extend to Gen-AI apps, while our categories capture the additional emphasis on clarifying AI behavior and managing expectations around evolving AI capabilities.
Developer response behavior differs across app categories, with some (i.e Education) prioritizing fast, template-based replies and others (i.e Games) slower but more substantive engagement. Developer response behavior also varies substantially across app categories as shown in Table 9. For example, Games exhibits both the highest reply ratio (56%) and the longest response delay (44.7 days). Our manual analysis suggests that this delay is associated with responses that confirm issues have been fixed, which represent 30% of replies in this category compared to only 2% in the overall dataset. Developers often appear to wait until a patch is deployed before responding, resulting in slower but more substantive communication.
@@
In contrast, Education apps demonstrate the shortest response delays (0.7 days). In our manual sample, 73% of responses fall under Acknowledgment and Appreciation, suggesting that developers frequently rely on quick, template-based replies that require minimal investigation. Education apps also show a relatively high proportion of Request for Additional Information responses (16%), reflecting situations where users report incorrect AI answers without providing sufficient context for developers to reproduce and resolve the issue.
@@
Meanwhile, Entertainment apps show the lowest reply ratio (2%). Developers in this category rarely respond to routine or positive feedback and instead appear to focus their replies on actionable problem reports. A large proportion of replies fall under Request for Additional Information (38%), Guidance and Troubleshooting Tips (22%), and Acknowledged and Under Investigation (19%), suggesting that developers primarily intervene when users report crashes, playback failures, or account issues that require additional diagnostic detail.
Developer responsiveness varies across Gen-AI topics, prioritizing actionable issues over systemic and human-centered concerns (Table 10). The highest reply ratio appears in Accessibility & Inclusivity (47%), despite this category containing the smallest number of reviews. This suggests that developers treat accessibility-related feedback as high-impact issues requiring immediate attention. Higher engagement also appears in Utility & Use Cases (18%) and Content Quality (14%), both of which directly affect user satisfaction and day-to-day usage of Gen-AI functionality.
@@
Emotional Connection 25,142 542 2% 7.5
@@
In contrast, topics such as Customer Support & Community (1%), Emotional Connection (2%), and Content Policy & Censorship (5%) receive fewer developer replies. These issues often involve systemic behavior of AI models or platform policies that cannot be easily modified, which may explain the lower engagement levels. Overall, this pattern suggests that developers prioritize feedback that can be addressed through short-term fixes (e.g., accessibility barriers or incorrect outputs), while responding less frequently to concerns that require deeper model changes or policy revisions.